# LLM Assertiveness can be Mechanistically Decomposed into Emotional and Logical Components

### v1

@@ 2.1 Datasets and model fine-tuning (Phase I-II)
We fine-tuned a Llama-3.2-1B-Instruct model (Grattafiori et al., 2024) with Low-Rank Adaptation (LoRA (Hu et al., 2021)) to predict the assertiveness scores. Datasets were separated into training and validation sets in each fold, one entire dataset (e.g. GM) served as validation ($\approx$ 130 items) and the remaining as training ($\approx$ 515 items) for five folds. Model performance was reported as mean squared error (MSE). After 30 epochs, the mean minimal MSE across the five folds reached 1.94, reducing from 61.11 at the first epoch.
@@ 2.2 Activation preprocessing and mechanistic interpretability (Phase III-IV)
To probe internal structure, we projected activations of the high-assertive items into a t-SNE map. Spatial groupings were examined to reveal data-driven sub-components (e.g., certainty vs persuasion) for later conceptual interpretation. Steering vectors from each sub-component were selectively removed, and Root Mean Squared Error (RMSE) changes in assertive prediction were measured to test whether these representations function independently or interact.

### v2

@@ 2.1 Datasets and model fine-tuning (Phase I-II)
We fine-tuned a Llama-3.2-1B-Instruct model (Grattafiori et al., 2024) with Low-Rank Adaptation (LoRA (Hu et al., 2021)) to predict the assertiveness scores. Datasets were separated into training and validation sets in each fold, one entire dataset (e.g. GM) served as validation ($\approx$ 130 items) and the remaining as training ($\approx$ 515 items) for five folds. Model performance was reported as mean squared error (MSE). After 30 epochs, the mean minimal MSE across the five folds reached 1.94, reducing from 61.11 at the first epoch (See Appendix A.1.1 for details).
@@ 2.2 Activation preprocessing and mechanistic interpretability (Phase III-IV)
To probe internal structure, we projected activations of the high-assertive items into a t-SNE map. Spatial groupings were examined to reveal data-driven sub-components (e.g., certainty vs persuasion) for later conceptual interpretation. Steering vectors from each sub-component were selectively removed, and Root Mean Squared Error (RMSE) changes in assertive prediction were measured to test whether these representations function independently or interact (see Appendix A.1.2 for details).
@@ 4 Conclusion and limitations
5 Acknowledgements
This work was partially completed during the SPAR (Supervised Program for Alignment Research) programme. We are especially grateful to Kellin Pelrine (Mila/FAR.AI), Dr. Maximilian Puelma Touzel (Mila/University of Montreal), Prof. Jean-François Godbout (Mila/University of Montreal), and Bijean Ghafouri (University of Southern California) for their mentorship, guidance, and valuable feedback throughout this project.